OpenStudy (anonymous):

When an altitude is drawn to the hypotenuse of a right triangle, the lengths of the segments of the hypotenuse are 9 and 36. What is the length of the altitude? I think it's 18

OpenStudy (anonymous):

I think you can use similar triangles to find\[{9 \over h} = {h \over 36}\]\[h^2 = 9 \times 36 = 9 \times 9 \times 4 = (9 \times 2)^2 = 18^2\]h=18 is the correct answer.